A solution is prepared by mixing 44.0 g of acetone (C3H6O) and 297.2g of chloroform (CHCl3). The vapor pressures of pure acetone and pure chloroform at 35 degrees C are 345 and 293 torr. If this is an ideal solution, what is the vapor pressure, in torr, of the solution at 35 degrees C?
Answer #1

Molar mass of acetone is 58 g/mole

Molar mass of chloroform is 119.5 g/mole

Number of moles of acetone , n = mass/molar mass

                                              = 0.76 moles

Number of moles of chloroform , n' = mass/molar mass

                                                 = 2.49 moles

Total number of moles , N = n + n' = 3.25 moles

Mole fraction of acetone , X = n/N = 0.234

We know that from Raoult's law (Po-P)/Po = X

Where

Po = vapor pressure of pure chloroform = 293 torr

P = vapor pressure of solution = ?

Plug the values we get (293-P)/293 = 0.234

                                    293-P = 68.5

                                                P = 224.5 torr